# Learning Shortcuts: On the Misleading Promise of NLU in Language Models

### 17 Jan 2024 v1

@@ 1 Introduction
Unfortunately, LLMs have been shown to exploit spurious associations and dataset biases as ‘shortcuts’, achieving inflated scores on NLU benchmarks, while lacking reliability and generalization on out-of-distribution samples. Recent works have shown that LLMs tend to learn shortcuts based on statistical cues e.g. the word “not” [24], keywords [22, 10] and cues related to language variations [23] to make predictions. This behavior, also known as shortcut learning, leads the model to learn non-generalizable decision rules that do not perform well on out-of-distribution data [22, 9], but continues to give state-of-the-art results on IID samples. Thus, PLM performance gains on NLU benchmarks do not necessarily portray improvement in the semantic and reasoning capabilities of language models.
This paper delineates the challenges in enhancing NLU capabilities within Pre-trained Language Models amidst the presence of shortcut learning. We offer a concise survey of recent research endeavors on quantifying and mitigating shortcut learning, examining its implications for language understanding in language models. Finally, we underscore the research gaps and leverage them to chart guidelines and future directions for advancing the field.
@@ 2.1 Identifying Shortcuts
While most works do not differentiate between spurious and genuine tokens, [27] addresses this distinction. The authors employ a human attention-guided approach to detect and alleviate shortcut learning. This mechanism identifies both model and dataset biases, utilizing human attention as a supervisory signal to compel the model to allocate more attention to ’relevant’ tokens. Their approach enhances robustness on out-of-distribution data and maintains performance on in-distribution (IID) data.
@@ 3.1.1 Inflated NLU Performance Scores
Moreover, [21] established that training and evaluating models using standard natural language inference (NLI) datasets such as MNLI [37] result in models primarily relying on heuristics. When assessed on a curated dataset lacking such heuristics, model performance dropped to $10\%$ for examples without heuristics, falling below chance performance. This highlights that NLI models may not inherently learn the correct rules for inference. Similarly, [29] shows that even with a complete vocabulary overlap, the performance of fact verification models declined by $20\%$ when employing distinct language—devoid of shortcuts—while maintaining consistent claim and evidence relations (support or refute). [27] also revealed a substantial decline of $40\%-60\%$ in BERT-base model performance on Natural Language Inference (NLI) and fact verification tasks upon the removal of shortcuts. These studies underscore the pronounced dependence of LLMs on superficial information within the datasets.
@@ 3.2.1 Data Centric Approaches
Recent approaches to enhance model robustness and mitigate the impact of spurious cues involve the creation of datasets that contain curated examples that reduce model reliance on shortcuts during task learning. Incorporating adversarial data points during both training and evaluation, such as mirroring linguistic artifacts around labels has been shown to diminish the impact of spurious cues on model learning [24, 29]. [29] demonstrated a notable decrease in model performance when assessed on an adversarial test set, underscoring the necessity for Natural Language Understanding (NLU) evaluations to incorporate considerations for shortcut learning in Language Models (LMs). Additional efforts in this domain involve the development of adversarial datasets, such as HANS, which assesses whether Natural Language Inference (NLI) models rely on syntactic heuristics to accomplish the task [21].
More recently, novel data generation techniques have been proposed for mitigating spurious correlations and promoting generalization [38]. This technique involves training data generators to create and filter data points that exacerbate spurious correlations but yield significant performance drops in task accuracy. Moreover, even though data augmentation for bias removal reduces the model’s dependence on spurious cues, it does not completely eliminate it [40]. This persistence of lexical data biases emphasizes the ongoing challenges in effectively mitigating the impact of spurious cues in NLP models.
@@ 4.1 Language and Vocabulary Variation
Accurately identifying and eliminating specific shortcuts used by Language Models (LLMs) poses a significant challenge. This is exemplified by the reliance on the word ”not” in BERT, where the removal of such shortcuts results in a notable drop in model performance. Successfully addressing this challenge requires nuanced strategies for pinpointing and mitigating these specific linguistic cues. Addressing the impact of shortcuts becomes complex when variations in language and vocabulary are introduced. Despite incorporating long-range context, encoding intricate lexical semantic phenomena for improved natural language understanding remains an ongoing endeavor [11, 1, 34]. Even when the vocabulary overlap is significant, maintaining consistent performance poses a challenge. This raises questions regarding the broader applicability of current state-of-the-art LLMs across different language contexts and the necessity for models to exhibit resilience and adaptability in the face of varied linguistic expressions.
@@ 4.2 Out-of-Distribution Generalization
Mitigating the impact of shortcuts becomes particularly challenging when Language Models (LLMs) are assessed on out-of-distribution datasets. The reliance on heuristics in natural language inference (NLI) models, particularly when trained and evaluated on standard NLI datasets like MNLI [37], poses a challenge to achieving robust inference capabilities. This is highlighted by a significant drop in performance when assessed on curated datasets lacking such heuristics, portraying difficulty in achieving generalization beyond the training distribution. Strategies and datasets need to be devised for the existing variety of NLU tasks, to enhance model robustness and adaptability to diverse data distributions.
@@ 5 Conclusion
This paper underscores the interplay between Natural Language Understanding (NLU) evaluation and Shortcut Learning (SL) within language models. By shedding light on the potential distortions SL introduces to NLU evaluations, we emphasize the critical need for a nuanced understanding of these impacts. It is necessary to actively address the challenges posed by SL, both in terms of identifying and mitigating shortcuts and refining evaluation methodologies. Neglecting these issues jeopardizes the reliability and fairness of NLU assessments. As we move forward, collaborative efforts within the NLP community are essential to advance our comprehension of SL, fostering the development of more robust language models and elevating the standards of NLU evaluation in real-world contexts.

### 09 Feb 2024 v2

@@ 1 Introduction
Unfortunately, LLMs have been shown to exploit spurious associations and dataset biases as ‘shortcuts’, achieving inflated scores on NLU benchmarks, while lacking reliability and generalization on out-of-distribution samples. Recent works have shown that LLMs tend to learn shortcuts based on statistical cues e.g. the word “not” [24], keywords [22, 10] and cues related to language variations [23] to make predictions. This behavior, also known as shortcut learning, leads the model to learn non-generalizable decision rules that do not perform well on out-of-distribution data [22, 9], but continues to give state-of-the-art results on independent identically distributed (IID) samples. Thus, PLM performance gains on NLU benchmarks do not necessarily portray improvement in the semantic and reasoning capabilities of language models.
This paper delineates the challenges in enhancing NLU capabilities within pre-trained language models amidst the presence of shortcut learning. We survey recent research on quantifying and mitigating shortcut learning and examine their implications for language under-standing in language models. Finally, we suggest potential research directions and guidelines that can help develop robust and reliable language models.
@@ 2.1 Identifying Shortcuts
While most works do not differentiate between spurious and genuine tokens, [27] addresses this distinction. The authors employ a human attention-guided approach to detect and alleviate shortcut learning. This mechanism identifies both model and dataset biases, utilizing human attention as a supervisory signal to compel the model to allocate more attention to ’relevant’ tokens. Their approach enhances robustness on out-of-distribution data and maintains performance on IID data.
@@ 3.1.1 Inflated NLU Performance Scores
Moreover, [21] established that training and evaluating models using standard NLI datasets such as MNLI [37] result in models primarily relying on heuristics. When assessed on a curated dataset lacking such heuristics, model performance dropped to $10\%$ for examples without heuristics, falling below chance performance. This highlights that NLI models may not inherently learn the correct rules for inference. Similarly, [29] shows that even with a complete vocabulary overlap, the performance of fact verification models declined by $20\%$ when employing distinct language—devoid of shortcuts—while maintaining consistent claim and evidence relations (support or refute). [27] also revealed a substantial decline of $40\%-60\%$ in BERT-base model performance on NLI and fact verification tasks upon the removal of shortcuts. These studies underscore the pronounced dependence of LLMs on superficial information within the datasets.
@@ 3.2.1 Data Centric Approaches
Recent approaches to enhance model robustness and mitigate the impact of spurious cues involve the creation of datasets that contain curated examples that reduce model reliance on shortcuts during task learning. Incorporating adversarial data points during both training and evaluation, such as mirroring linguistic artifacts around labels has been shown to diminish the impact of spurious cues on model learning [24, 29]. [29] demonstrated a notable decrease in model performance when assessed on an adversarial test set, underscoring the necessity for Natural Language Understanding (NLU) evaluations to incorporate considerations for shortcut learning in LLMs. Additional efforts in this domain involve the development of adversarial datasets, such as HANS, which assesses whether NLI models rely on syntactic heuristics to accomplish the task [21].
More recently, novel data generation techniques have been proposed for mitigating spurious correlations and promoting generalization [38]. This technique involves training data generators to create and filter data points that exacerbate spurious correlations but yield significant performance drops in task accuracy. Moreover, even though data augmentation for bias removal reduces the model’s dependence on spurious cues, it does not eliminate it [40]. This persistence of lexical data biases emphasizes the ongoing challenges in effectively mitigating the impact of spurious cues in NLP models.
@@ 4.1 Language and Vocabulary Variation
Accurately identifying and eliminating specific shortcuts used by LLMs poses a significant challenge. This is exemplified by the reliance on the word ”not” in BERT, where the removal of such shortcuts results in a notable drop in model performance. Successfully addressing this challenge requires nuanced strategies for pinpointing and mitigating these specific linguistic cues. Addressing the impact of shortcuts becomes complex when variations in language and vocabulary are introduced. Despite incorporating long-range context, encoding intricate lexical semantic phenomena for improved natural language understanding remains an ongoing endeavor [11, 1, 34]. Even when the vocabulary overlap is significant, maintaining consistent performance poses a challenge. This raises questions regarding the broader applicability of current state-of-the-art LLMs across different language contexts and the necessity for models to exhibit resilience and adaptability in the face of varied linguistic expressions.
@@ 4.2 Out-of-Distribution Generalization
Mitigating the impact of shortcuts becomes particularly challenging when LLMs are assessed on out-of-distribution datasets. The reliance on heuristics in NLI models, particularly when trained and evaluated on standard NLI datasets like MNLI [37], poses a challenge to achieving robust inference capabilities. This is highlighted by a significant drop in performance when assessed on curated datasets lacking such heuristics, portraying difficulty in achieving generalization beyond the training distribution. Strategies and datasets need to be devised for the existing variety of NLU tasks, to enhance model robustness and adaptability to diverse data distributions.
@@ 5 Conclusion
This paper underscores the interplay between natural language understanding evaluation and shortcut learning within language models. We shed light on the distortions introduced by shortcut learning on NLU evaluation and emphasize the need for a nuanced understanding of its impacts. It is necessary to actively address the challenges posed by shortcut learning, both in terms of identifying and mitigating shortcuts and refining evaluation methodologies. Neglecting these issues jeopardizes the reliability and fairness of NLU assessments. As we move forward, collaborative efforts within the NLP community are essential to advance our comprehension of shortcut learning, fostering the development of more robust language models and elevating the standards of NLU evaluation in real-world contexts.